OpenStudy (anonymous):

Mayday Mayday! HELP! Given the equation for simple harmonic motion d=8 cos(3pi/4 t), find the period.

OpenStudy (anonymous):

the equation for simple harmonic motion has the following form\[d = A \cos \left( \omega t \right)\]we can identify that\[A = 8\]\[\omega = \frac{3 \pi}{4}\]the angular frequency is\[\omega = \frac{2 \pi}{T}\]\[\frac{3 \pi}{4} = \frac{2\pi}{T}\]solve for T, T is the time period

OpenStudy (anonymous):

 What's your answer?